OpenStudy (anonymous):

A country's population in 1993 was 72million. In 1996 it was 76 million. Estimate the population in 2003 using the exponential growth formula. Round your answer to the nearest million.

OpenStudy (anonymous):

did you get it right?

OpenStudy (anonymous):

1993 is year 0 1996 is year 3 2003 is year 10

OpenStudy (anonymous):

compute \[72\times( \frac{76}{72})^{\frac{10}{3}}\]

OpenStudy (anonymous):

we can use formula if you like. i mean exponential growth formula

OpenStudy (anonymous):

I keep getting those wrong :( Idk why :( The bacteria ones i right in like 10 seconds.

OpenStudy (anonymous):

good! must be something about adjusting those years. check your mistakes to see what you are doing wrong, and you will stop

OpenStudy (anonymous):

I got 86.218 How do i round it to the nearest million?

OpenStudy (anonymous):

86 million!

OpenStudy (anonymous):

oops D:

OpenStudy (anonymous):

btw your math teacher will want you to do something different. i can walk you through that if you like